8.      Mr. Toriop owns 5000 shares of stock in Yummy Corporation. The company has announced that it will pay a dividend of $5 per share in one year and then a liquidating dividend of $50 per share in two years. The required return on ABC stock is 10%. (21 marks)

a. What is the current share price of your stock?                           (1 mark)

b. What will be the company’s share price in one year?                   (1 mark)

c. Mr. Toriop wishes to have equal amount of dividend income for the next two years. How can he use homemade leverage on Yummy Corporation’s dividends to achieve this goal? Check that the present value of the cash flows will be the same as they are before the homemade leverage. (Hint: Dividends will be in the form of an annuity.) (8 marks)

d. Suppose Mr. Toriop is thinking about buying a house for $220,000 in one year. How can he use homemade leverage on Yummy Corporation’s dividends to achieve this goal? Check that the present value of the cash flows will be the same as they are before the homemade leverage.                               (5 marks)

e. Suppose Mr. Toriop is thinking about postponing the house purchase for two years, by which time the price of the house will have increased by $35,000. How can he use homemade leverage on Yummy Corporation’s dividends to achieve this goal? Check that the present value of the cash flows will be the same as they are before the homemade leverage.      (6 marks)
Answer #1
No. of shares Given 5000
Dividend Given 5
Liquidating dividend Given 50
Required return Given 10%
Discount factor Formula 1/(1+Discount rate)^No. of years
Year Cash flow Discount factor Discounted cash flow
1 5 0.9091 4.55
2 50 0.8264 41.32
Current stock price 45.87
Company's share price in 1 year
Year Cash flow Discount factor Discounted cash flow
0 5 1.0000 5.00
1 50 0.9091 45.45
Price after 1 year 50.45
Price after 2 years is same as value of liquidating dividend i.e. 50.
Dividend after 1 year 5000*5 25000
Price after 1 year 50.45
Price after 2 years 50
Since cash flow of both years to be equal
Let no. of shares to be sold after 1 year to generate homemade dividend be x.
25000+(x*50.45) = (5000-x)*50
25000+50.45x = 250000-50x
100.45x = 225000
x= 225000/100.45
x= 2239.920358
Therefore, sell 2240 shares after receiving 1st year's dividend to get homemade dividend.
Cash flow after one year
From sale of shares 2240*50.45 113008
From dividend 5000*5 25000
138008
Cash flow after 2 years
From liquidating dividend (5000-2240)*50 139242
If cost of house is 220,000
Cash flow from dividend 5000*5 25000
Balance cost to be recovered from sale of shares 220000-25000 195000
Price per share 50.45
No. of shares to be sold 195000/50.45 3,865.21
By selling 3866 shares, the goal can be achieved.
If cost of house increases to 220000+35000= 255000
Cash flow of first year 5000*5 25000
Liquidating dividend of 2nd year 5000*50 250000
275000
There is no need to utilize homemade dividends.
The cash flow arising at the end of 2 years will be sufficient to cover the cost of house.